# Non-abelian Hodge correspondence over singular Kähler spaces

### v1

@@ 5. Non-abelian Hodge correspondence over Kähler klt spaces
This section first proves the two key ingredients outlined in Section 1.3: Theorem 1.12 (in Section 5.1) and Theorem 1.15 (in Section 5.2). We then prove Theorems 1.3 and 1.7 in Section 5.3, thereby establishing the nonabelian Hodge correspondences for compact Kähler klt spaces. Finally, we prove Corollary 1.9, Corollary 1.10 and Theorem 1.11 in Section 5.4.
@@
The proof of Theorem 1.12 will be divided into two parts. We review the following statements, which are useful in the proof of Theorem 1.12 and 1.15.
@@
Let $0\rightarrow\mathcal{S}\rightarrow\mathcal{E}\rightarrow\mathcal{Q}\rightarrow 0$ be a short exact sequence of coherent sheaves on a Cohen-Macaulay, normal and irreducible complex space $X$ of dimension $n$. Suppose that $\mathcal{E}$ is reflexive, $\mathcal{S}$ and $\mathcal{Q}^{\vee\vee}$ is locally free. If addition $\mathcal{Q}$ is locally free in codimension $2$, then $\mathcal{E}$ and $\mathcal{Q}$ are locally free.
Lemma 5.1 follows from the following result that come back to Scheja and Trautmann [95, 103].
Proposition 5.2 (cf. [100, Pages 142]).
Let $X$ be a complex space, $A\subset X$ a closed analytic subset and $\mathcal{F}$ be a coherent sheaf on $X$. Fix a non-negative integer $k$. Then the following statements are equivalent:
@@
    for every open subset $U\subset X$, the restrictions $H^{i}(U,\mathcal{F})\rightarrow H^{i}(U\setminus A,\mathcal{F})$ are bijective for $i<k$ and injective for $i=k$.
We give an explanation on (Proposition 5.2 $\Rightarrow$ Lemma 5.1) for reader’s convenience. By the assumption of Lemma 5.1, there exists an analytic subspace $Z\subset X$ of codimension at least $3$ such that $\mathcal{Q}$ and $\mathcal{E}$ is locally free outside $Z$. Since $\mathcal{Q}^{\vee}=\mathcal{Q}^{\vee\vee\vee}$ and $\mathcal{S}$ are locally free, then for any $x\in X$,
@@
where $\mathrm{codh}$ denotes the depth and third equality follows from the dimension formula. Thus the singularity set $S_{n-1}(\mathcal{S}\otimes\mathcal{Q}^{\vee}):=\{x\in X:\mathrm{codh}(\mathcal{S}\otimes\mathcal{Q}^{\vee})_{x}\leq n-1\}$ of $\mathcal{S}\otimes\mathcal{Q}^{\vee}$ is empty and $S_{n}(\mathcal{S}\otimes\mathcal{Q}^{\vee})=X$, we have that $\dim(A\cap S_{\mu+3})\leq\mu\ \text{ for any integer }\mu.$ Therefore, the restriction $H^{1}(X,\mathcal{S}\otimes\mathcal{Q}^{\vee})\rightarrow H^{1}(X\setminus Z,\mathcal{S}\otimes\mathcal{Q}^{\vee})$ is bijective by Proposition 5.2. This implies $\mathcal{F}|_{X\setminus Z}$ extends to a locally free Higgs sheaf $\mathcal{F}^{\prime}$ on $X$ that is the extension of $\mathcal{S}$ by $\mathcal{Q}^{\vee\vee}$. By the reflexivity, $\mathcal{F}$ coincides with $\mathcal{F}^{\prime}$ and the proof is complete.
Lemma 5.3 (The Hodge index Theorem, cf. [53, Lemma 3.17]).
@@
The equality holds if and only if $\widehat{c}_{1}(\mathcal{E})\cdot[\gamma]\cdot[\alpha_{0}]\cdots[\alpha_{n-2}]=0$ for any $\gamma\in H^{2}(X,\mathbb{R})$.
@@ 5.1.1. Characterization of semistable reflexive Higgs sheaf by extension
Let $X$ be a compact Kähler klt space of dimension $n$, $\gamma:Y\rightarrow X$ be a maximally quasi-étale cover of $X$ and $(\mathcal{E}_{X_{\mathrm{reg}}},\theta_{\mathcal{E}_{X_{\mathrm{reg}}}})$ be a reflexive Higgs sheaf over $X_{\mathrm{reg}}$. Then the following statements are equivalent:
@@
    $\mathcal{E}_{Y}=\gamma^{[*]}\mathcal{E}_{X}$ is locally free, in particular $\gamma^{*}\theta_{\mathcal{E}_{X_{\mathrm{reg}}}}$ can be extended to a Higgs field $\theta_{Y}$ of $\mathcal{E}_{Y}$; $\mathcal{E}_{Y}$ admits a filtration of $\theta_{Y}$-invariant locally free sheaves
@@
    such that each quotient $\mathcal{Q}_{k}:=\mathcal{E}_{k}/\mathcal{E}_{k-1}$ is locally free and its all Chern classes vanish; $\mathcal{Q}_{k}$ with the induced Higgs field $\theta_{k}$ is $(\gamma^{*}\omega_{0},\cdots,\gamma^{*}\omega_{n-2})$-stable with respect to any Kähler forms $\omega_{0},\cdots,\omega_{n-2}$ on $X$.
@@
    Orbifold Chern classes $\widehat{c}_{1}(\mathcal{E}_{X}),\widehat{c}_{1}^{2}(\mathcal{E}_{X}),\widehat{c}_{2}(\mathcal{E}_{X})$ of $\mathcal{E}_{X}$ vanish and $(\mathcal{E}_{X_{\mathrm{reg}}},\theta_{\mathcal{E}_{X_{\mathrm{reg}}}})$ is semistable with respect to any Kähler forms $\omega_{0},\cdots,\omega_{n-2}$ on $X$.
Ditto for polystability (in this case, the length $l$ of the filtration equals $1$).
@@
The implication $(iii)\Rightarrow(i)$ is obvious and $(ii)\Rightarrow(iii)$ follows from Lemma 2.9, Lemma 2.24 and Lemma 2.21. It suffices to verify $(i)\Rightarrow(ii)$. According to the same reason of provining $(ii)\Rightarrow(iii)$, we may assume that $X$ is maximally quasi-étale and $\gamma$ is the identity map.
If $(\mathcal{E}_{X_{\mathrm{reg}}},\theta_{\mathcal{E}_{X_{\mathrm{reg}}}})$ is $(\omega_{0},\cdots,\omega_{n-2})$-stable, then the proof is completed by applying Theorem 1.13, Proposition 4.14 and Lemma 4.1. Because any reflexive Higgs sheaf of $\operatorname{rank}1$ is $(\omega_{0},\dots,\omega_{n-2})$-stable, we could argue by induction on $\operatorname{rank}(\mathcal{E}_{X})$. Assuming that Theorem 5.4 holds for $\operatorname{rank}(\mathcal{E}_{X})<r$. Now let $\operatorname{rank}(\mathcal{E}_{X})=r$. Without loss of generality, we assume that $(\mathcal{E}_{X{\mathrm{reg}}},\theta_{\mathcal{E}_{X{\mathrm{reg}}}})$ is not $(\omega_{0},\dots,\omega_{n-2})$-stable.
The construction of $\mathcal{F}_{X}$ in this paragraph is standard, we include an explanation for reader’s convenience. It follows from Remark 2.6 that
@@
is nonempty. Taking $\mathcal{F}_{X}\in\mathcal{B}$ with the maximal rank. We may assume that $\mathcal{F}_{X}$ is saturated by Remark 2.6. In particular, the induced Higgs sheaf $(\mathcal{F}_{X_{\mathrm{reg}}},\theta_{\mathcal{F}_{X_{\mathrm{reg}}}})$ is $(\omega_{0},\cdots,\omega_{n-2})$-semistable. Set $\mathcal{Q}_{X_{\mathrm{reg}}}:=\mathcal{E}_{X_{\mathrm{reg}}}/\mathcal{F}_{X_{\mathrm{reg}}}$. Then a direct computation gives
@@
The induced torsion-free Higgs sheaf $(\mathcal{Q}_{X_{\mathrm{reg}}},\theta_{\mathcal{S}_{X_{\mathrm{reg}}}})$ must be $(\omega_{0},\cdots,\omega_{n-2})$-stable. Otherwise, there exists a saturated subsheaf $\mathcal{G}_{X}\subsetneq\mathcal{Q}_{X}$ that is $\theta_{\mathcal{Q}_{X_{\mathrm{reg}}}}$-invariant such that $\mu_{(\omega_{0},\cdots,\omega_{n-2})}(\mathcal{G}_{X})\geq\mu_{(\omega_{0},\cdots,\omega_{n-2})}(\mathcal{Q}_{X}).$ Let $\mathcal{S}_{X}=\mathcal{G}_{X}+\mathcal{F}_{X}$, then we have the short exact sequence
$0\rightarrow\mathcal{G}_{X}\rightarrow\mathcal{S}_{X}\rightarrow\mathcal{F}_{X}\rightarrow 0,$
$\mathcal{S}_{X_{\mathrm{reg}}}$ is $\theta_{\mathcal{E}_{X_{\mathrm{reg}}}}$-invariant and we have $\mu_{(\omega_{0},\cdots,\omega_{n-2})}(\mathcal{S}_{X})\geq\mu_{(\omega_{0},\cdots,\omega_{n-2})}(\mathcal{E}_{X})$ by a direct computation, which contradicts that $\mathcal{F}_{X}$ has the maximal rank in $\mathcal{B}$.
It follows from Proposition 2.15, Bogomolov-Gieseker inequality (3.1) and Lemma 5.3 that
(5.2) $\begin{split}0=&\widehat{\mathrm{ch}}_{2}(\mathcal{E}_{X})\cdot[\omega_{1}]\cdots[\omega_{n-2}]\\ \leq&\widehat{\mathrm{ch}}_{2}(\mathcal{F}_{X})\cdot[\omega_{1}]\cdots[\omega_{n-2}]+\widehat{\mathrm{ch}}_{2}(\mathcal{Q}_{X}^{\vee\vee})\cdot[\omega_{1}]\cdots[\omega_{n-2}]\\ \leq&\frac{1}{\operatorname{rank}(\mathcal{F}_{X})}\widehat{c}_{1}^{2}(\mathcal{F}_{X})\cdot[\omega_{1}]\cdots[\omega_{n-2}]+\frac{1}{\operatorname{rank}(\mathcal{Q}_{X})}\widehat{c}_{1}^{2}(\mathcal{Q}^{\vee\vee})\cdot[\omega_{1}]\cdots[\omega_{n-2}]\\ \leq&\frac{1}{\operatorname{rank}\mathcal{F}_{X}}\frac{\big(\widehat{c}_{1}(\mathcal{F}_{X})\cdot[\omega_{0}]\cdots[\omega_{n-2}]\big)^{2}}{[\omega_{0}]\cdot[\omega_{0}]\cdots[\omega_{n-2}]}+\frac{1}{\operatorname{rank}\mathcal{F}_{X}}\frac{\big(\widehat{c}_{1}(\mathcal{Q}_{X}^{\vee\vee})\cdot[\omega_{0}]\cdots[\omega_{n-2}]\big)^{2}}{[\omega_{0}]\cdot[\omega_{0}]\cdots[\omega_{n-2}]}=0\end{split}$
since $\mu_{(\omega_{0},\cdots,\omega_{n-2})}(\mathcal{E}_{X})=0$. Hence, each inequality is in fact an equality, which implies the following statements:
@@
    $(\mathcal{F}_{X_{\mathrm{reg}}},\theta_{\mathcal{F}_{X_{\mathrm{reg}}}})$ is a $(\omega_{0},\cdots,\omega_{n-2})$-semistable reflexive Higgs sheaf with (1.3); By the induction hypothesis, $\mathcal{F}_{X}$ is locally free and $(\mathcal{F}_{X_{\mathrm{reg}}},\theta_{\mathcal{F}_{X_{\mathrm{reg}}}})$ satisfies property (2) of Theorem 5.4.
@@
    $(\mathcal{Q}_{X_{\mathrm{reg}}},\theta_{\mathcal{Q}_{X_{\mathrm{reg}}}})^{\vee\vee}$ is a $(\omega_{0},\cdots,\omega_{n-2})$-stable reflexive Higgs sheaf with (1.3); Building on the stable case that has been clarified, $\mathcal{Q}_{X}^{\vee\vee}$ is locally free and its all Chern classes vanish.
@@
    $\mathcal{Q}_{X}^{\vee\vee}=\mathcal{Q}_{X}$ in codimension 22 (see Proposition 2.15); Buiding on Lemma 5.1, $\mathcal{Q}_{X}$ and $\mathcal{E}_{X}$ is locally free. In particular, the Higgs fields $\theta_{\mathcal{E}_{X_{\mathrm{reg}}}}$ and $\theta_{\mathcal{Q}_{X_{\mathrm{reg}}}}$ extend to the Higgs fields on $\mathcal{E}_{X}$ and $\mathcal{Q}_{X}$, respectively.
Summarizing the above statements and the induction hypothesis, the proof is complete. ∎
5.1.2. Independent of the choice of the polarization
Let us complete the remaining of Theorem 1.12 by proving the following.

### v2

@@ 5. Non-abelian Hodge correspondence on Kähler klt spaces
This section first proves the two key ingredients outlined in Section 1.3: Theorem 1.12 (in Section 5.1) and Theorem 1.15 (in Section 5.2). We then prove Theorems 1.3 and 1.7 in Section 5.3, thereby establishing the nonabelian Hodge correspondences for compact Kähler klt spaces. Finally, we prove Corollaries 1.9 and 1.10 and Theorem 1.11 in Section 5.4.
@@
The proof of Theorem 1.12 will be divided into two parts. We review the following statements, which are useful in the proof of Theorems 1.12 and 1.15.
@@
Let $0\rightarrow\mathcal{S}\rightarrow\mathcal{E}\rightarrow\mathcal{Q}\rightarrow 0$ be a short exact sequence of coherent sheaves on a Cohen-Macaulay, normal and irreducible complex space $X$ of dimension $n$. Suppose that $\mathcal{E}$ is reflexive, $\mathcal{S}$ and $\mathcal{Q}^{\vee\vee}$ are locally free. If, in addition, $\mathcal{Q}$ is locally free in codimension $2$, then both $\mathcal{E}$ and $\mathcal{Q}$ are locally free.
Lemma 5.1 follows from a result of Scheja and Trautmann [98, 106].
Proposition 5.2 (cf. [103, Page 142]).
Let $X$ be a complex space, $A\subset X$ a closed analytic subset and $\mathcal{F}$ a coherent sheaf on $X$. Fix a nonnegative integer $k$. Then the following statements are equivalent:
@@
    for every open subset $U\subset X$, the restriction maps $H^{i}(U,\mathcal{F})\rightarrow H^{i}(U\setminus A,\mathcal{F})$ are bijective for $i<k$ and injective for $i=k$.
For the reader’s convenience, we sketch how Proposition 5.2 implies Lemma 5.1. By the assumptions of Lemma 5.1, there exists an analytic subspace $Z\subset X$ of codimension at least $3$ such that $\mathcal{Q}$ and $\mathcal{E}$ are locally free outside $Z$. Since $\mathcal{Q}^{\vee}=\mathcal{Q}^{\vee\vee\vee}$ and $\mathcal{S}$ are locally free, for any $x\in X$, we have
@@
where $\mathrm{codh}$ denotes the depth, and the third equality follows from the dimension formula. Thus the singularity set $S_{n-1}(\mathcal{S}\otimes\mathcal{Q}^{\vee}):=\{x\in X:\mathrm{codh}(\mathcal{S}\otimes\mathcal{Q}^{\vee})_{x}\leq n-1\}$ of $\mathcal{S}\otimes\mathcal{Q}^{\vee}$ is empty, and $S_{n}(\mathcal{S}\otimes\mathcal{Q}^{\vee})=X$. Hence $\dim(A\cap S_{q+3})\leq q\ \text{ for any integer }q.$ Therefore, the restriction $H^{1}(X,\mathcal{S}\otimes\mathcal{Q}^{\vee})\rightarrow H^{1}(X\setminus Z,\mathcal{S}\otimes\mathcal{Q}^{\vee})$ is bijective by Proposition 5.2. This means that $\mathcal{E}|_{X\setminus Z}$ extends to a locally free sheaf $\mathcal{E}^{\prime}$ on $X$ that is the extension of $\mathcal{S}$ by $\mathcal{Q}^{\vee\vee}$. By reflexivity of $\mathcal{E}$, we have $\mathcal{E}$ coincides with $\mathcal{E}^{\prime}$. Hence $\mathcal{E}$ and $\mathcal{Q}$ are locally free. This completes the proof.
Lemma 5.3 (Hodge Index Theorem, cf. [54, Lemma 3.17]).
@@
The equality holds if and only if $\widehat{c}_{1}(\mathcal{Q})\cdot[\sigma]\cdot[\alpha_{1}]\cdots[\alpha_{n-2}]=0$ for every $\sigma\in H^{2}(X,\mathbb{R})$.
@@ 5.1.1. Characterization of semistable reflexive Higgs sheaf by extension
Let $X$ be a compact Kähler klt space of dimension $n$, $\gamma:Y\rightarrow X$ be a maximally quasi-étale cover and $(\mathcal{E}_{X_{\mathrm{reg}}},\theta_{\mathcal{E}_{X_{\mathrm{reg}}}})$ be a reflexive Higgs sheaf over $X_{\mathrm{reg}}$. Then the following statements are equivalent:
@@
    The pullback $\mathcal{E}_{Y}:=\gamma^{[*]}\mathcal{E}_{X}$ is locally free, and $\gamma^{*}\theta_{\mathcal{E}_{X_{\mathrm{reg}}}}$ extends to a Higgs field $\theta_{Y}$ on $\mathcal{E}_{Y}$. Moreover, $\mathcal{E}_{Y}$ admits a filtration by $\theta_{Y}$-invariant locally free subsheaves
@@
    such that each quotient $\mathcal{Q}_{k}:=\mathcal{E}_{k}/\mathcal{E}_{k-1}$ is locally free with all Chern classes vanishing, and the induced locally free Higgs sheaf $(\mathcal{Q}_{k},\theta_{k})$ is $(\gamma^{*}\omega_{0},\cdots,\gamma^{*}\omega_{n-2})$-stable with respect to any Kähler forms $\omega_{0},\cdots,\omega_{n-2}$ on $X$.
@@
    The orbifold Chern classes $\widehat{c}_{1}(\mathcal{E}_{X}),\widehat{c}_{1}^{2}(\mathcal{E}_{X}),\widehat{c}_{2}(\mathcal{E}_{X})$ vanish, and $(\mathcal{E}_{X_{\mathrm{reg}}},\theta_{\mathcal{E}_{X_{\mathrm{reg}}}})$ is semistable with respect to any Kähler forms $\omega_{0},\cdots,\omega_{n-2}$ on $X$.
The same holds for polystability, in which case the filtration in (ii) has length $l=1$.
@@
The implication $(iii)\Rightarrow(i)$ is obvious, and $(ii)\Rightarrow(iii)$ follows from Lemma 2.9, Lemma 2.24 and Proposition 2.21. Thus it suffices to verify $(i)\Rightarrow(ii)$. By the same reasoning as in the implication $(ii)\Rightarrow(iii)$, we may assume that $X$ itself is maximally quasi-étale, so that $\gamma$ is the identity.
If $(\mathcal{E}_{X_{\mathrm{reg}}},\theta_{\mathcal{E}_{X_{\mathrm{reg}}}})$ is $(\omega_{0},\cdots,\omega_{n-2})$-stable, the conclusion follows directly from Theorem 1.13, Proposition 4.14 and Lemma 4.1. Note that any reflexive Higgs sheaf of $\operatorname{rank}1$ is $(\omega_{0},\dots,\omega_{n-2})$-stable. We will argue by induction on $\operatorname{rank}(\mathcal{E}_{X})$. Assume that Theorem 5.4 holds for all ranks smaller than $r$, and let $\operatorname{rank}(\mathcal{E}_{X})=r$. Without loss of generality, we assume that $(\mathcal{E}_{X{\mathrm{reg}}},\theta_{\mathcal{E}_{X{\mathrm{reg}}}})$ is not $(\omega_{0},\dots,\omega_{n-2})$-stable.
The following construction of $\mathcal{F}_{X}$ is standard; we include the details for the reader’s convenience. It follows from Remark 2.6 that the set
@@
is nonempty. Choose $\mathcal{F}_{X}\in\mathcal{B}$ of maximal rank. We may assume that $\mathcal{F}_{X}$ is saturated by Remark 2.6. In particular, the induced Higgs sheaf $(\mathcal{F}_{X_{\mathrm{reg}}},\theta_{\mathcal{F}_{X_{\mathrm{reg}}}})$ is $(\omega_{0},\cdots,\omega_{n-2})$-semistable. Set $\mathcal{Q}_{X_{\mathrm{reg}}}:=\mathcal{E}_{X_{\mathrm{reg}}}/\mathcal{F}_{X_{\mathrm{reg}}}$. Then a direct computation gives
@@
We claim that the induced torsion-free Higgs sheaf $(\mathcal{Q}_{X_{\mathrm{reg}}},\theta_{\mathcal{Q}_{X_{\mathrm{reg}}}})$ is $(\omega_{0},\cdots,\omega_{n-2})$-stable. Indeed, if it were not, there would exist a $\theta_{\mathcal{Q}_{X_{\mathrm{reg}}}}$-invariant saturated subsheaf $\mathcal{G}_{X}\subsetneq\mathcal{Q}_{X}$ with $\mu_{(\omega_{0},\cdots,\omega_{n-2})}(\mathcal{G}_{X})\geq\mu_{(\omega_{0},\cdots,\omega_{n-2})}(\mathcal{Q}_{X}).$ Consider a proper subsheaf $\mathcal{S}_{X}\subsetneq\mathcal{E}_{X}$ such that $\mathcal{S}_{X}/\mathcal{F}_{X}=\mathcal{G}_{X}$. Then we have a short exact sequence
$0\rightarrow\mathcal{F}_{X}\rightarrow\mathcal{S}_{X}\rightarrow\mathcal{G}_{X}\rightarrow 0.$
The sheaf $\mathcal{S}_{X_{\mathrm{reg}}}$ is $\theta_{\mathcal{E}_{X_{\mathrm{reg}}}}$-invariant, and a direct computation shows $\mu_{(\omega_{0},\cdots,\omega_{n-2})}(\mathcal{S}_{X})\geq\mu_{(\omega_{0},\cdots,\omega_{n-2})}(\mathcal{E}_{X})$, contradicting the maximality of $\operatorname{rank}(\mathcal{F}_{X})$ in $\mathcal{B}$.
Now, applying Proposition 2.15, the Bogomolov-Gieseker inequality (3.1), and Lemma 5.3, we obtain that
(5.2) $\begin{split}0=&\widehat{\mathrm{ch}}_{2}(\mathcal{E}_{X})\cdot[\omega_{1}]\cdots[\omega_{n-2}]\\ \leq&\widehat{\mathrm{ch}}_{2}(\mathcal{F}_{X})\cdot[\omega_{1}]\cdots[\omega_{n-2}]+\widehat{\mathrm{ch}}_{2}(\mathcal{Q}_{X}^{\vee\vee})\cdot[\omega_{1}]\cdots[\omega_{n-2}]\\ \leq&\frac{1}{\operatorname{rank}(\mathcal{F}_{X})}\widehat{c}_{1}^{2}(\mathcal{F}_{X})\cdot[\omega_{1}]\cdots[\omega_{n-2}]+\frac{1}{\operatorname{rank}(\mathcal{Q}_{X})}\widehat{c}_{1}^{2}(\mathcal{Q}^{\vee\vee})\cdot[\omega_{1}]\cdots[\omega_{n-2}]\\ \leq&\frac{1}{\operatorname{rank}\mathcal{F}_{X}}\frac{\big(\widehat{c}_{1}(\mathcal{F}_{X})\cdot[\omega_{0}]\cdots[\omega_{n-2}]\big)^{2}}{[\omega_{0}]^{2}\cdot[\omega_{1}]\cdots[\omega_{n-2}]}+\frac{1}{\operatorname{rank}\mathcal{Q}_{X}}\frac{\big(\widehat{c}_{1}(\mathcal{Q}_{X}^{\vee\vee})\cdot[\omega_{0}]\cdots[\omega_{n-2}]\big)^{2}}{[\omega_{0}]^{2}\cdot[\omega_{1}]\cdots[\omega_{n-2}]}=0\end{split}$
since $\mu_{(\omega_{0},\cdots,\omega_{n-2})}(\mathcal{F}_{X})=\mu_{(\omega_{0},\cdots,\omega_{n-2})}(\mathcal{Q}_{X})=0$. Hence, every inequality in (5.2) is an equality. This yields the following three conclusions:
@@
    $(\mathcal{F}_{X_{\mathrm{reg}}},\theta_{\mathcal{F}_{X_{\mathrm{reg}}}})$ is an $(\omega_{0},\cdots,\omega_{n-2})$-semistable reflexive Higgs sheaf satisfying (1.3). By the induction hypothesis, $(\mathcal{F}_{X_{\mathrm{reg}}},\theta_{\mathcal{F}_{X_{\mathrm{reg}}}})$ satisfies (ii) of Theorem 5.4. In particular, $\mathcal{F}_{X}$ is locally free.
@@
    $(\mathcal{Q}_{X_{\mathrm{reg}}},\theta_{\mathcal{Q}_{X_{\mathrm{reg}}}})^{\vee\vee}$ is an $(\omega_{0},\cdots,\omega_{n-2})$-stable reflexive Higgs sheaf satisfying (1.3); Consequently, by the stable case already established, $\mathcal{Q}_{X}^{\vee\vee}$ is locally free and all its Chern classes vanish.
@@
    $\mathcal{Q}_{X}^{\vee\vee}=\mathcal{Q}_{X}$ in codimension 22 (see Proposition 2.15); Using Lemma 5.1, we conclude both $\mathcal{Q}_{X}$ and $\mathcal{E}_{X}$ are locally free. In particular, the Higgs fields $\theta_{\mathcal{E}_{X_{\mathrm{reg}}}}$ and $\theta_{\mathcal{Q}_{X_{\mathrm{reg}}}}$ extend to Higgs fields on $\mathcal{E}_{X}$ and $\mathcal{Q}_{X}$, respectively.
Putting together the above statements and the induction hypothesis completes the proof. ∎
5.1.2. Independence of the choice of Kähler polarization
We now complete the proof of Theorem 1.12 by establishing the following result.